# Intergroup violence in bursts or fizzles

### v10

@@ Abstract
During intergroup confrontations, agitating stimuli such as opponents’ threats and provocations can trigger collective violence, even when the usual mechanisms of ingroup cooperation, such as norms with sanctions, are absent. We examine video recordings of street fights between groups of young men. Their violence sometimes breaks out in a burst, while at other times it reaches only a fizzle in which only a few group members participate. An adapted Ising model explains the dilemma of collective action and demonstrates that these two temporal unfoldings can be predicted by the proportion of defectors in a focal group.
@@ 1 Introduction
In studies of intergroup conflict, the large numbers of casualties of wars between and within countries draw most attention [17, 59]. Yet, the most frequent occurrences of intergroup violence involve small groups, including subgroups of larger groups. We focus on small groups, and ask how violence committed by non-professionals unfolds. When observing pertinent videos, one notices a burst or a fizzle in the majority of cases. Before punches are thrown, individuals in the focal group and their opponents insult and threaten, and thereby agitate, one another. For this to result in violence by focal group members, they have to be in close proximity and have to have face-to-face contact [19, 16, 45].
In one of the most influential studies of violence, Randall Collins [19] argues that antagonists need to circumvent an emotional barrier of tension and fear in order to be able to use violence. Surmounting the barrier requires that assailants attain a state of emotional dominance. In face-to-face interpersonal conflicts, they can reach emotional dominance when the other party appears vulnerable or weaker, or when the focal group is supported by an audience [19]. Collins’ notion of “forward panics” corresponds to the bursts we observed. In forward panics, assailants seize the opportunity to forcefully attack opponents in moments of weakness or vulnerability, for example when being isolated or falling down to the ground. This emotional barrier raises another barrier known as the dilemma of collective action [53, 54], which has gained less attention in scholarly work on violence. We will present a formal model that explicates the conditions for groups to burst into collective violent action. At the core of the model is opponents’ agitation.
When agitation passes a critical level (for example, when one individual is pushed by an opponent), violence can break out in a burst wherein a majority of individuals (or a subgroup from a larger and more dispersed group) participates. Alternatively, the confrontation can fizzle out, even if group members are agitated. In the case of fizzling, one or few individuals push or punch asynchronously, but there is no burst of violence. Additionally, understanding how collective violence unfolds is complicated by other people present on the scene, who may form an audience or try to de-escalate [39, 57, 79]. De-escalation happens often in street fights [58]. Given a common prelude of proximity and agitation mixed with de-escalation, our question is how to explain bursts and fizzles. To this end we use the centennial Ising model from physics [43, 70]. It resulted in a Nobel Prize for Parisi (in 2021; [55]), and has been used in various studies of social influence [14, 69, 34, 26]. We adapt the model in a novel way, use it as an agent based model, and apply it to videos of street fights between groups of (mostly) young men.
@@ 2 Theory
Collective violence involves a dilemma of collective action [54, 53], so before predicting bursts and fizzles, we must first address how the Ising model explains overcoming such dilemmas. The face-to-face contacts of the focal group members are the social ties of their network. We already know from public goods experiments that if participants regard a public good (in our case the attack of or defense against opponents) to be valuable, more than half of the participants are conditional cooperators willing to contribute if others do [15]. In other words, they conform to their (weighted) average neighbor in the network [7, 73]. Our empirical study is about relatively inexperienced individuals, rather than seasoned fighters, hence our subjects experience uncertainty when they come under threat, which increases the likelihood of their conforming to the group [81, 49]. Conformity makes sense from an evolutionary perspective when payoffs are hard to predict [75]. Therefore, subjects’ motivation not only depends on their appraisal of the public good, but also their desire to align with their group members. [^1] The interesting question of why certain people self-select into groups where they can fight against others is beyond the scope of this study. Milliff [47] suggested that these people perceive their situation as uncertain but controllable, which is consistent with our assumptions.
Now think of two individuals, each with the behavioral options to defect or cooperate. Behavioral options correspond to magnetic spins in the original Ising model, but that is not relevant here; we apply the model differently and do not establish an analogy with the original. The possibilities are: (1) both individuals defect, which avoids exploitation but does not yield any public good; (2) one individual freerides the cooperating other, yielding half of the public good; or (3) both cooperate at a cost, which maximizes the public good. The dashed line in Fig. 1 plots these possibilities, with the normalized number of cooperators ($N_{C}/n$) from left to right. At the beginning, both defect (the local minimum of $H$ on the left). The vertical axis ($H$, or energy in the original) conveys that moving uphill (i.e., that one of the two will start contributing to the public good) is unlikely. After all, most individuals understand that fighting alone against a group is very dangerous. By contrast, if moving to lower levels of $H$ is possible, it will certainly happen; thus, if one of the two cooperates, the other conditional cooperator will join in, shifting $H$ to its global minimum on the right. The hill is the graphical representation of the dilemma of collective action; it is also drawn for a group of five. It shows that without a motivating force, it is very unlikely that a group of non-fighting individuals ($n\geq 2$) will overcome the hill and fight collectively, and it is even less likely in larger groups.
@@
Nonetheless, some individuals may start fighting, which in turn may prompt others to join. When agitating stimuli, such as opponents’ threats, increase, focal group members become aroused and angry. Consequently, one or few individuals may accidentally cooperate even when most others do not (yet). This phenomenon of accidental, or spontaneous, cooperation—called “trembling hands” in game theory [21]—can influence proximate others and may result in a cascade of cooperation.
@@
For the model, we define turmoil ($T$, or temperature in the original) as opponents’ behavior that agitates members of the focal group, which (in our case) means that their arousal and anger increase; the hormones adrenaline and cortisol are produced and linger on for a while. In the model, opponents’ turmoil has an effect on the focal group at the aggregate level and affects everyone equally. In larger groups than ours, individuals standing closer to opponents are exposed to more turmoil than others standing far away. Locally varying turmoil in the Ising model is discussed elsewhere [10]; here, we avoid unnecessary complications. When keeping it simple, the model demonstrates that if turmoil increases, there is no gradual increase of accidental cooperators. Instead, when turmoil reaches a critical level ($T_{c}$), fighting breaks out in a burst; see Figure 2. Our contribution is to show that the outbreak of collective violence depends on a second critical level, namely the critical proportion of unconditional defectors in the focal group ($p_{c}$). In violent situations, many people do not fight, for all kinds of reasons. They may be too scared to fight [19], have empathy with their opponents, disagree with violence (i.e., appraise the public good differently), feel no solidarity with their group [18], try to de-escalate, be stopped in their tracks by de-escalators (also called guardians), or they may have fought but were wounded, were wrestled to the ground, or became exhausted at some point and became passive. If their proportion stays below the critical level ($p\leq p_{c}$), turmoil ($T>T_{c}$) will be followed by a burst of violence. Above the critical level ($p>p_{c}$), however, there is neither a critical level of turmoil nor a burst, and violence occurs in a fizzle (Fig. 2). The critical level of unconditional defectors does not follow from earlier theories, but the Ising model allows us to infer it, which we use as a prediction that we test empirically. Note that $p$ incorporates not only de-escalators in the focal group, but also focal group members who have been prevented from fighting by de-escalators in the focal group or by others.
@@
Figure 2: Color online. Simulation of Equation 1 over a range of turmoil ($T$) on the horizontal axis. The proportion of cooperators ($N_{C}/n$) in a fully connected network in the size range of our empirical study ($n=5$) is on the vertical axis; $C=1$ and $D=1/2$. The black line (top) depicts a burst in the group without unconditional defectors ($p=0$), the red line (middle) a burst in the group with one unconditional defector ($p<p_{c}$), and the dashed blue line (bottom) a fizzle in the group with two unconditional defectors ($p>p_{c}$).
@@
Most models of collective action in the literature revolve around individual rewards and punishments, called selective incentives [53], on top of a share of the public good. These incentives require norms about (in)appropriate behavior in a given situation [23], as well as monitoring of group members [61], and transmission of information (i.e., gossip) through the group’s network, which leads to reputations [51] that feedback through selective incentives, with or without leaders. This package of mechanisms is crucial for ongoing cooperation in the longer run but needs time to develop, which may not be available when threats are imminent. The more time people have, the better they can prepare themselves, which is especially important for high-risk situations. Examples of well-prepared groups are police, soldiers, firefighters, and combat medics who receive professional training that enables them to cooperate effectively and respond to situational stimuli in predetermined manners rather than spontaneously. By contrast, ordinary citizens who face violent opponents are far less prepared, or not prepared at all. For non-professional groups, such as ours, the uncertainties (of outcomes, benefits, and costs) are higher.
@@
Several earlier models of cooperation do not feature the usual package of cooperation mechanisms; namely, models of thresholds [30], cascades [77], and critical mass [44]. Therefore, these models are potentially useful for our study. They draw on the assumption that some people take the initiative or leaders set cooperation in motion. It has been found that in protests, however, leaders often proscribe violence, and most violence occurs in the absence of leaders [32]. Hence, the presence of leaders does not explain the outbreak of violence in general. The Ising model, in contrast, points out how cooperation can start spontaneously without leaders, galvanized by accidental cooperators rather than exceptionally zealous ones. If there are initiative takers or leaders in favor of violence [28], however, they can be accommodated in our model, and people may also enter and leave a focal group.
@@
These models also draw on strong rationality assumptions, such as perfect information on the numbers of cooperators at every moment. Yet, the degree of rationality is bounded in violent and other uncertain situations due to incomplete information about opponents and group members beyond face-to-face contact. Moreover, inexperienced fighters cannot assess their payoffs with any degree of accuracy before a fight. The Ising model is more parsimonious than the aforementioned models because it does not rely on assumptions of strong rationality. Individuals may have fairly accurate expectations of their payoffs, but for the model it is enough if they find the public good valuable or not. The model combines the precision of game theory, including a mapping on payoffs—unknown to the participants—with the behavioral spontaneity of chaotic and uncertain situations [71], which is usually assessed qualitatively [68], as well as a portion of randomness in individuals’ decisions, reflecting their uncertainty [42]. Moreover, whereas game theoretic approaches to conflict readily become complicated, even for two individuals [64], the simplicity of the Ising model makes it possible to investigate large networks computationally.
@@
Sociology’s most influential, qualitative model of violence has been developed by Collins [19]. Our work contributes to his model. One of its important contributions was that it shifted researchers’ attention to what actually happens in fights, for which they started using visual data (e.g., [50]). At the core, his model has potentially quantifiable variables capturing emotions, such as confrontational tension (agitation in our model), forward panic (our burst), and emotional dominance, to predict the beginning of violence and the winner in the end. Empirical studies using these variables have remained qualitative [8, 31, 36], to which we add different, quantitative insights. We build on Collins’ work, also use visual data, but our general strategy is using behavioral variables that are relatively easily observable. We minimize the number of variables from the literature that are hard to observe behaviorally, such as reputations, norms, emotions, and solidarity, without downplaying their importance; they can be incorporated in an extended version of the Ising model [9] but we lack video recorded data to include them in our study. Complementing Collins, our approach is explicitly relational, adds precision, and explains collective violent action. We will first demonstrate how the Ising model explains surmounting collective action’s barrier.
@@ 3 Model
Members of a (fledgling) focal group, indexed $i$ or $j$, can defect, $D$, or contribute, $C$, to a public good, with $0<D<C$. Behavioral variable $S_{i}$ of conditional cooperators can take the value $S_{i}=C$ or $S_{i}=-D$, whereas unconditional defectors stay put at $S_{j}=-D$. Before a collective action, everyone in a focal group defects. Network ties among focal group members, $A_{ij}>0$, mean that $i$ is in close proximity to, and senses the behavior of, group member $j$, or else $A_{ij}=0$. Because people tend to respond to proportions of their social environment rather than absolute numbers [77, 25], ties are row-normalized, with $w_{ij}=A_{ij}/\sum_{j=1}^{n}A_{ij}$ such that $\sum_{j=1}^{n}w_{ij}=1$. Based on our video data, we assume that in out small groups ($n<10$), sensing is reciprocal at least to some degree, but not necessarily symmetrical, and that everyone senses all others, thus the network is fully connected. [^2] Temporary exceptions to full connectedness in our data were groups where some people’s view was blocked by objects, opponents, or bystanders, as well as one larger group ($n=14$). These cases we investigated through simulations. The model can be used for complex networks as well, beyond the small groups studied here [10].
The Ising model is the following Hamiltonian equation [5, 76]
@@
We do not assume that individuals know their payoffs in advance, but they will heuristically—and perhaps wrongly—distinguish between valuable ($C>D$) and unvaluable ($C<D$) public goods or, to the same effect, between potentially efficacious and non-efficacious actions [62]. Note that payoffs are not used in the model’s calculations but are defined to provide a meaningful interpretation. When $i$ chooses between $C$ and $D$ amid $N_{C}$ cooperators, payoffs for cooperation and defection are, respectively
@@
For our empirical study, we have to choose values for $C$ and $D$ to predict the critical level of unconditional defectors, $p_{c}$. The most obvious choice is $C=1$, as in game theory and the original Ising model. For $D$ we want to avoid two trivial values: if $D=1$, there is no point in cooperating, which is equally valuable as defecting, and if $D=0$, there is no dilemma (but a downward slope to the right in Fig. 1) and violence becomes self-explanatory. Choosing $D$ to be maximally distant from the two trivial values seems to be a reasonable first approximation. Hence, we set $S=\{1,-1/2\}$ for all conditional cooperators, also in the examples. These two values were published in a theoretical study preceding this empirical study [11]. For unconditional defectors we set $S_{j}=-1/2$, irrespective of their reasons, simply because they all discourage the conditional cooperators to fight (Fig. 1). Unconditional means under the conditions in our cases; the notion juxtaposes them to conditional cooperators. Note that all earlier Ising models had either the values $\{1,-1\}$ [14, 69, 26, 80] or $\{1,0\}$ [20]. Of all these models, only one represents a public goods game, in this case for two individuals [1, 63] whereas our model is applicable to groups of any size.
[…]
@@
The computational approach, over a certain range of $T$, is as follows (Fig. 3). For a given level of turmoil, a network node $i$ is randomly picked and $H$ is calculated. For comparison, $i$’s current behavior is flipped, from $D$ to $C$ (or the other way around if $i$ cooperates), and $H^{\prime}$ with the flip is calculated. The flip is accepted and implemented if $H^{\prime}<H$ or with a certain chance that increases with $T$. In other words, increasing $T$ increases the amount of randomness (due to situational turmoil) in $i$’s decision. A behavioral change of $i$ affects others in $i$’s neighborhood when it is their turn to decide. Consecutive decisions are Monte Carlo steps in the Metropolis algorithm [5] that loops through great many Monte Carlo steps (counted by $t$ in Fig. 3) in order to allow individuals’ interdependent behavior to settle down. This procedure is repeated at increments of $T$, in our study with step size 0.01.
@@
Figure 3: The Metropolis algorithm. Index $i$ runs over the conditional cooperators, not the unconditional defectors; $t$ is a counter of Monte Carlo steps; $c_{r}$ is a random number in the range $0\leq c_{r}\leq 1$.
@@
To illustrate the emergence of cooperation with increasing turmoil in a group of $n=5$, we apply the Metropolis algorithm; see Figure 2. First, there are only conditional cooperators. At low turmoil, collective action does not start (black line at the top), but at a critical level $T_{c}$, (almost) everybody bursts into cooperation, with a maximum (where $N_{C}/n\approx 1$) at or near $T_{c}$. Then, a small number of accidental cooperators win over most others to join the collective action. The effect of turmoil is nonmonotonic and the level of cooperation decreases if $T$ keeps increasing beyond $T_{c}$, which means that extraordinarily strong turmoil becomes more confusing than motivating to fight.
@@
If there are unconditional defectors, $T_{c}$ increases, which is more pronounced in larger networks than in Figure 2, and maximum cooperation is lower than in the previous case because the number of conditional cooperators is lower. [^5] In Fig. 1, a proportion of unconditional defectors, $p$, means that no matter how many conditional cooperators contribute, $N_{C}/n\leq 1-p$. If the proportion of unconditional defectors surpasses a critical level, $p_{c}$, and $T$ increases above $T_{c}$, there is a fizzle (dashed blue line in Fig. 2) with much lower levels of cooperation, and no burst-like start of cooperation. In the mean field analysis, there is no (burst of) cooperation if $p_{c}>S_{0}/\Delta$, independent of network size and density. In simulations, $p_{c}$ is proportionally less precise in smaller networks (Table S3), but because our networks are complete, we stick to the mean field prediction. Based on the mean field analysis ($p_{c}=1/3$) and a nearly identical result from the simulations ($p_{c}=0.34$), [^6] If unconditional defectors are clustered together in a sparse network, they are less in the way of collective action of the remainder network (thus $p_{c}$ is higher) than if they are evenly spread out across the network. our main prediction is that $p_{c}=1/3$.

### v12

@@ Abstract
During intergroup confrontations, agitating stimuli such as opponents’ threats and provocations can trigger collective violence, even without the usual mechanisms of ingroup cooperation, such as norms with sanctions. We examine video recordings of street fights between groups of young men. Their violence sometimes breaks out in a burst, wherein a majority of participants starts fighting almost simultaneously. At other times, only few group members participate and it takes them more time to do so. This difference in commencing collective violence can be understood by perceiving it as a collective action dilemma. We adapt an Ising model to show that the proportion of group members who cannot or do not want to contribute to the public good—victory over opponents—predicts whether violence takes the form of a burst or not.
@@ 1 Introduction
In studies of intergroup conflict, the large numbers of casualties of wars between and within countries draw most attention [18, 64]. Yet, the most frequent occurrences of intergroup violence involve small groups, including subgroups of larger groups. We focus on small groups, and want to explain how violence committed by unarmed, non-professionals starts off. When observing pertinent videos, one often notices bursty starts of violence, where the majority of a group engages almost simultaneously in violence against their opponents. Before punches are thrown, individuals in the focal group and their opponents insult and threaten, and thereby agitate, one another. For this to result in collective violence by focal group members, a critical level of agitation has to be reached. A necessary condition is that focal group members are in close proximity and maintain face-to-face contact [20, 17, 49].
In one of the most influential studies of violence, Randall Collins [20] argues that antagonists need to circumvent an emotional barrier of tension and fear in order to be able to use violence. Surmounting the barrier requires that assailants attain a state of emotional dominance. In face-to-face interpersonal conflicts, they can reach emotional dominance when the other party appears vulnerable or weaker, or when the focal group is supported by an audience [20]. Collins’ notion of “forward panics” corresponds to the bursts we observed. In forward panics, assailants seize the opportunity to forcefully attack opponents in moments of weakness or vulnerability, for example when being isolated or falling down to the ground. This emotional barrier is meshed with another barrier known as the dilemma of collective action [58, 59], which has gained less attention in scholarly work on violence. We will present a formal model that explicates the conditions for groups to burst into collective violent action. At the core of the model is opponents’ agitation.
When agitation passes a critical level (for example, when one individual is pushed by an opponent), violence can break out in a burst. Alternatively, the confrontation can escalate more slowly and with fewer participants. Understanding how collective violence starts is complicated by the presence of other people, who may form an audience or try to de-escalate [42, 62, 84]. De-escalation happens often in street fights [63]. Given the conditions of proximity, agitation, and de-escalation, our goal is to explain bursts versus non-bursts. To this end we use the centennial Ising model from physics [46, 75]. It resulted in a Nobel Prize for Parisi (in 2021; [60]) and has been used in various studies of social influence [15, 74, 37, 29]. We adapt the model in a novel way, use it as an agent based model, and apply it to videos of street fights between groups of (mostly) young men.
@@ 2 Theory
Collective violence involves a dilemma of collective action [59, 58], so before predicting bursts, we must first address how the Ising model explains overcoming such dilemmas. For an unarmed focal group, the ingroup ties that matter are face-to-face contacts in close proximity [20]. We already know from public goods experiments that if participants regard a public good to be valuable, more than half of the participants are conditional cooperators willing to contribute if many others do [16]. In other words, conditional cooperators conform to their (weighted) average neighbor in the network [7, 78]. In our case, the public good is victory over opponents, which means that opponents flee or are wrestled to the ground. If victory is not feasible, a second tier public good is the prevention of ingroup humiliation. When confronted with opponents, most individuals experience uncertainty, which increases the likelihood of their conforming to the group [86, 53]. Conformity makes sense from an evolutionary perspective when payoffs are hard to predict [80]. Therefore, individuals’ motivation not only depends on their appraisal of the public good, but also their desire to align with their group members.
Now think of two individuals, each with the behavioral options to defect or cooperate. Behavioral options correspond to magnetic spins in the original Ising model, but that is not relevant here; we apply the model differently and do not lean on an analogy with the original. The possibilities are: (1) both individuals defect, which avoids exploitation but does not yield any public good; (2) one individual cooperates, yielding half of the public good, but is exploited by the freeriding other; or (3) both cooperate at a cost, which maximizes the public good. The dashed line in Fig. 1 plots these possibilities, with the normalized number of cooperators ($N_{C}/n$) from left to right. At the beginning, both defect (the local minimum of $H$ on the left). The vertical axis ($H$, or energy in the original) conveys that moving uphill (i.e., that one of the two will start contributing to the public good) is unlikely. After all, most individuals understand that fighting alone against a group is very dangerous. By contrast, if moving to lower levels of $H$ is possible, it will certainly happen; thus, if one of the two cooperates, the other conditional cooperator will join in, shifting $H$ to its global minimum on the right. The hill is the graphical representation of the dilemma of collective action; it is also drawn for a group of five. It shows that without a motivating force, it is very unlikely that a group of non-fighting individuals ($n\geq 2$) will overcome the hill and fight collectively, and collective action is even less likely in larger groups.
@@
Nonetheless, some individuals may start fighting, which in turn may prompt others to join. When agitating stimuli, such as opponents’ threats, increase, focal group members become aroused and angry. Consequently, one or few individuals may accidentally cooperate even when most others do not (yet). This phenomenon of accidental or spontaneous cooperation is called “trembling hands” in game theory [23]. Among defectors, a few trembling hands can influence proximate others and may initiate a cascade of cooperation, or burst if the cascade proceeds rapidly. Among cooperators, trembling hands can have the opposite effect.
For the model, we define turmoil ($T$, or temperature in the original) as opponents’ behavior that agitates members of the focal group. Agitation arouses them; they produce the hormones adrenaline and cortisol that linger on for a while. In the model, opponents’ turmoil has an effect on the focal group at the aggregate level and affects everyone equally. In larger groups than ours, individuals standing closer to opponents are exposed to more turmoil than others standing further away. Locally varying turmoil in the Ising model as well as network variations are examined elsewhere [10]; here, we make the model not more complicated than necessary for our empirical study. The model demonstrates that if turmoil increases, there is no gradual increase of accidental cooperators. Instead, when turmoil reaches a critical level, $T_{c}$, fighting breaks out in a burst; see Figure 2. Bursts clearly differ from a gradual build up of collective action and the distinction does not follow from other models in the literature. One might expect, by contrast, that the tipping point is smooth rather than sudden (it is a second order phase transition; [55]), and in large random networks (e.g., $n=1000$) it is. However, all large social networks are decomposed into smaller clusters, where the transition is sudden rather than smooth, also in time (Fig. 5), and in small clusters, the tipping point is reached first [10].
@@
The distinction between bursty and slowly increasing collective action turns out to depend on the critical proportion of unconditional defectors (versus conditional cooperators) in the focal group, $p_{c}$. Perhaps they would be conditional cooperators under different circumstances, but they were not in our study. Hence, we use $p_{c}$ as a prediction that we test empirically. In violent situations, many people do not fight, for all kinds of reasons. They may be too scared to fight [20], have empathy with their opponents, disagree with violence (i.e., appraise the public good differently), feel no solidarity with their group [19], try to de-escalate, be stopped in their tracks by de-escalators (also called guardians in criminological studies), or they may have fought but were wounded, were wrestled to the ground, or became exhausted at some point and became passive. In the population at large, preferences for violence will have a distribution [26] with a negative mean value that expresses a general dislike of violence. Only few individuals in the tail of the distribution will self-select into violent groups and value the public good positively. [^1] Milliff [51] suggested that these people perceive their situation as uncertain but controllable. If the proportion of unconditional defectors in these groups stays below the critical level ($p\leq p_{c}$), turmoil ($T>T_{c}$) will be followed by a burst of violence. Above the critical level ($p>p_{c}$), however, there is neither a critical level of turmoil nor a burst (Fig. 2). Note that $p$ incorporates not only de-escalators in the focal group, but also focal group members who have been prevented from fighting by de-escalators in the focal group or by others.
@@
Figure 2: Simulation of Equation 1 over a range of turmoil ($T$) on the horizontal axis. The proportion of cooperators ($N_{C}/n$) in a fully connected network in the size range of our empirical study ($n=5$) is on the vertical axis; $C=1$ and $D=1/2$. The black line (top) depicts a burst in the group without unconditional defectors ($p=0$). The dashed blue line (bottom) depicts a gradual increase of cooperation in the group with two unconditional defectors ($p>p_{c}$).
@@
Most models of collective action in the literature revolve around individual rewards and punishments, called selective incentives [58], on top of a share of the public good. These incentives require norms about (in)appropriate behavior in a given situation [25], as well as monitoring of group members [66], and transmission of information (i.e., gossip) through the group’s network, which leads to reputations [56] that feedback through selective incentives, with or without leaders. This package of mechanisms is crucial for ongoing cooperation in the longer run but needs time to develop, which may not be available when threats are imminent. The more time people have, the better they can prepare themselves, which is especially important for high-risk situations. Examples of well-prepared groups are police, soldiers, firefighters, and combat medics who receive professional training that enables them to cooperate effectively and respond to situational stimuli in predetermined manners rather than spontaneously. By contrast, ordinary citizens and amateur fighters lack training and team coordination. For non-professional groups, such as ours, the uncertainties of outcomes, benefits, and costs are higher.
@@
Several earlier models of cooperation do not feature the usual package of cooperation mechanisms; namely, models of thresholds [33], cascades [82], and critical mass [48]. Therefore, these models are potentially useful for our study. They draw on the assumption that some people take the initiative or leaders set cooperation in motion. Yet in many cases, violence occurs in the absence of leaders [35]. Hence, an explanation of violence should not depend on their presence. The Ising model points out how cooperation can start spontaneously without leaders, galvanized by accidental cooperators rather than exceptionally zealous ones. If there are initiative takers or leaders in favor of violence [31], however, they can be accommodated in our model.
@@
These earlier models also draw on strong rationality assumptions, such as perfect information about the numbers of cooperators at every moment. Yet, in violent and other uncertain situations, the degree of rationality is bounded due to incomplete or absent information about opponents and group members beyond face-to-face contact. Moreover, individuals often respond to provocations in ways that harm their interests, by overestimating their own abilities and underestimating the tenacity of their opponents. The Ising model is more parsimonious than the aforementioned models because it does not rely on assumptions of strong rationality. For the model it is enough if individuals find the public good valuable, without assuming how (in)accurate their expectations are. The model combines the precision of game theory, including a mapping on payoffs—usually unknown to the participants—with the behavioral spontaneity of chaotic and uncertain situations [76]. Beyond threshold models, this uncertainty have been assessed qualitatively [73] instead of modeling it precisely [45]. Uncertainty has been added to threshold models, too, which makes it possible to loosen the rationality assumptions. Then, a small probability, $\epsilon$, that an individual starts cooperating by random chance suggests that the probability that someone in a group cooperates increases exponentially with group size, $1-(1-\epsilon)^{n}$ [47]. However, this argument omits the interdependence of group members. Consequently, it contradicts our empirical finding that cooperation starts earlier in smaller groups, as predicted by the Ising model. Finally, whereas game theoretic approaches to conflict readily become complicated, even if there are only two individuals [69], the simplicity of the Ising model makes it possible to investigate large networks computationally.
@@
Our work contributes to Collins’ [20] influential qualitative model of violence. One of its important contributions was that it shifted researchers’ attention to what actually happens in fights, for which they started using visual data. At the core, his model has potentially quantifiable variables capturing emotions, such as confrontational tension (agitation in our model), forward panic (our burst), and emotional dominance, to predict the beginning of violence and the dominating party. Empirical studies using these variables have remained qualitative [8, 34, 39], to which we contribute quantitative insights. We thus build on Collins’ work, also use visual data, but our general strategy is to use behavioral variables that are relatively easily observable. We minimize the number of variables from the literature that are hard to observe, such as reputations, norms, and solidarity, without downplaying their importance; they can be incorporated in an extended version of the Ising model [9] but we lack video recorded data to include them in our study. For example, if a pro-social norm is important and has been measured, it can be implemented as an external field (see below). Then, it will precipitate cooperation. Complementing Collins [20], our approach adds precision though modeling and explains collective action.
@@ 3 Model
Members of a (possibly fledgling) focal group, indexed $i$ or $j$, can defect, $D$, or contribute, $C$, to a public good, with $0<D<C$. Behavioral variable $S_{i}$ of conditional cooperators can take the value $S_{i}=C$ or $S_{i}=-D$, whereas unconditional defectors stay put at $S_{j}=-D$. Before a collective action, everyone in a focal group defects. Network ties among focal group members, $A_{ij}>0$, mean that $i$ is in close proximity to, and senses the behavior of, group member $j$, or else $A_{ij}=0$. Because people tend to respond to proportions of their social environment rather than absolute numbers [82, 28], ties are row-normalized, with $w_{ij}=A_{ij}/\sum_{j=1}^{n}A_{ij}$ such that $\sum_{j=1}^{n}w_{ij}=1$. Based on our video data, we assume that in our small groups ($n<10$), sensing is reciprocal at least to some degree, but not necessarily symmetrical, and that everyone senses all others, thus the network is fully connected. [^2] Temporary exceptions to full connectedness in our data were groups where some people’s view was blocked by objects, opponents, or bystanders, as well as one larger group ($n=14$). These cases we investigated through simulations. The Ising model is the following Hamiltonian equation [5, 81]
@@
We do not assume that individuals know their payoffs in advance, but they will heuristically—and perhaps wrongly—distinguish between valuable ($C>D$) and non-valuable ($C<D$) public goods or, to the same effect, between potentially efficacious and non-efficacious actions [67]. Note that payoffs are not used in the model’s calculations but are defined to provide a meaningful interpretation. When $i$ chooses between $C$ and $D$ amid $N_{C}$ cooperators, payoffs for cooperation and defection are, respectively
@@
For our empirical study, we have to choose values for $C$ and $D$ to predict the critical level of unconditional defectors, $p_{c}$. The most obvious choice is $C=1$, as in game theory and the original Ising model. For $D$ we want to avoid two trivial values: if $D=0$, there is no dilemma (but a downward slope to the right in Fig. 1), and if $D=1$, there is no point in cooperating, which is equally valuable as defecting. Choosing $D$ to be maximally distant from the two trivial values seems to be a reasonable first approximation. Hence, we set $S=\{1,-1/2\}$ for all conditional cooperators, also in the examples [11]. For unconditional defectors $j$ we set $S_{j}=-1/2$, irrespective of their reasons. Note that all earlier Ising models had either the values $\{1,-1\}$ [15, 74, 29, 85] or $\{1,0\}$ [21]. Of all these models, only one represents a public goods game, in this case for two individuals [1, 68] whereas our model is applicable to groups of any size.
Beyond our empirical study, the payoffs in the asymmetric Ising model can be generalized by relating $C$ and $D$ to the symmetric model through a mapping $\{C,-D\}\rightarrow\{S_{0}+\Delta,S_{0}-\Delta\}$, with a bias $S_{0}=(C-D)/2$ with respect to 0, and the two behavioral options symmetrical at each side of $S_{0}$ at an offset $\Delta=\pm(C+D)/2$. [^4] It can be shown that the asymmetry in $S$ is equivalent to the symmetric model with an external field $2S_{0}$ [11]. In the payoffs, the bias and offset are expressed through $R=S_{0}/\Delta$. If $\Delta$ is set to a fixed value (here, 0.75), decreasing $S_{0}$ (here, 0.25) makes the public good and cooperation for it less valuable. This decline is equivalent to an increasing proportion of cooperating network neighbors winning over an actor to cooperate, also in other binary decision models [82, 33]. Increasing $S_{0}$ makes cooperation more valuable, and it corresponds to a decreasing proportion of cooperating network neighbors. If there are initiative takers, indexed $j$, they will have a higher $S_{0,j}$ than the other group members, which lowers $T_{c}$ in small networks.
@@
The computational approach over a certain range of $T$ is as follows (Fig. 3). For a given level of turmoil, a network node $i$ is randomly picked and $H$ is calculated. For comparison, $i$’s current behavior is flipped, from $D$ to $C$ (or the other way around if $i$ cooperates), and $H^{\prime}$ with the flip is calculated. The flip is accepted and implemented if $H^{\prime}<H$ or with a certain chance that increases with $T$. In other words, increasing $T$ increases the amount of randomness (due to situational turmoil) in $i$’s decision. A behavioral change of $i$ affects others in $i$’s neighborhood when it is their turn to decide. Consecutive decisions are Monte Carlo steps in the Metropolis algorithm [5] that loops through great many Monte Carlo steps (counted by $t$ in Fig. 3) in order to allow individuals’ interdependent behavior to settle down. This procedure is repeated at increments of $T$, in our study with step size 0.01.
@@
Figure 3: The Metropolis algorithm. Index $i$ runs over the conditional cooperators, not the unconditional defectors; $t$ is a counter of Monte Carlo steps; and $c_{r}$ is a random number in the range $0\leq c_{r}\leq 1$.
@@
To illustrate the emergence of cooperation with increasing turmoil in a group of $n=5$, we apply the Metropolis algorithm; see Figure 2. First, we investigate the case without unconditional defectors. At low turmoil, collective action does not start (continuous line), but at a critical level $T_{c}$, (almost) everybody bursts into cooperation, with a maximum (where $N_{C}/n\approx 1$) close to $T_{c}$. Then, a small number of accidental cooperators wins over most others to join the collective action. The effect of turmoil is nonmonotonic and the level of cooperation decreases if $T$ keeps increasing beyond $T_{c}$, which means that extraordinarily strong turmoil becomes more confusing than motivating to fight.
@@
If there are unconditional defectors, $T_{c}$ increases, which is more pronounced in larger networks, and maximum cooperation is lower than in the previous case because the number of conditional cooperators is lower. [^5] In Fig. 1, a proportion of unconditional defectors, $p$, means that no matter how many conditional cooperators contribute, $N_{C}/n\leq 1-p$. If the proportion of unconditional defectors surpasses a critical level, $p_{c}$, and $T$ increases above $T_{c}$, there is no burst (dashed blue line in Fig. 2) and cooperation increases gradually to a low level. In the mean field analysis, there is no (burst of) cooperation if $p_{c}>S_{0}/\Delta$, independent of network size and density. In simulations, $p_{c}$ is proportionally less precise in smaller networks (Table S3), but because our networks are complete, we stick to the mean field prediction. Based on the mean field analysis ($p_{c}=1/3$) and a nearly identical result from the simulations ($p_{c}=0.34$), [^6] If unconditional defectors are clustered together in a sparse network, they are less in the way of collective action of the remainder network (thus $p_{c}$ is higher) than if they are evenly spread out across the network. our main prediction is that $p_{c}=1/3$.